# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with $L_{2}$-mean Hörmander conditions

### v1

@@ Abstract.
We construct a slightly new noncommutative Calderón-Zygmund decomposition by further splitting the bad function. Using this tool, we prove the weak type (1,1) boundedness of noncommutative Calderón-Zygmund operators under $L_{2}$-mean Hölder conditions, which improves the previous result.
@@ 1. Introduction
Calderón-Zygmund operators, a well-known class of singular integral operators, were introduced by Calderón and Zygmund [Calderon1952On] in their seminal 1952 work. The kernel of such an operator satisfies specific conditions, including the size and regularity conditions, the latter of which was refined by Hörmander [Hormander1960Estimates] in 1960.
@@
where $x$, $y$, $z\in\operatorname{\mathbb{R}}^{d}$. The Lipschitz regularity condition (1.2) can be relaxed to certain variants of smoothness conditions (see [Grafakos2019Alimited, Suzuki2021TheCalderon, Watson1990Weighted]). In the following, all Calderón-Zygmund kernels under consideration are of nonconvolution type.
Inspired by operator algebras, harmonic analysis, noncommutative geometry, and quantum probability (see [Cadilhac2018Weak, Chen2013Harmonic, Junge2014Smooth, Junge2018Noncommutative, Mei2007Operator, Mei2009Pseudo-localization, Mei2017Free, Mei2022Mikhlin, Parcet2009Pseudo-localization]), noncommutative harmonic analysis has become an exciting field in recent years. However, generalizing the Calderón-Zygmund decomposition, a key technique in harmonic analysis first introduced in [Calderon1952On], to the noncommutative setting is challenging.
In 2009, Parcet [Parcet2009Pseudo-localization] rigorously constructed a kind of noncommutative Calderón-Zygmund decomposition by using Cuculescu’s maximal weak type $(1,1)$ estimates for noncommutative martingales [Cuculescu1971Martingales]. With this tool, Parcet established the weak type $(1,1)$ estimates for operator-valued Calderón-Zygmund singular integrals with the kernel satisfying the Lipschitz regularity condition (1.2) (Cadilhac later gave a shorter proof in [Cadilhac2018Weak]). Notably, the Hörmander condition
@@
which is weaker than (1.2), is already sufficient for weak type $(1,1)$ boundedness in classical Calderón-Zygmund theory. It is therefore natural to ask whether the noncommutative weak type $(1,1)$ boundedness still holds under the Hörmander condition (1.3)? This question remains open.
A step forward was made in 2022 by Cadilhac et al. [Cadilhac2022Spectral]. They developed a more effective noncommutative Calderón-Zygmund decomposition without the off-diagonal term of the good function, which allowed them to prove the noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators if the kernel $K$ satisfies
@@
where $\ell(Q)$ and $c(Q)$ stand for the length and the center of the dyadic cube $Q$, respectively. More generally, for $1\leq q\leq\infty$, we say that a kernel $K$ satisfies the $L_{q}$-integral regularity condition if
@@
where $\delta_{q}(m)$ is defined as
$\sup_{{\begin{subarray}{c}y\in\operatorname{\mathbb{R}}^{d},\ R>0\\ |v|<R\end{subarray}}}$ $\Big((2^{m}R)^{d(q-1)}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x\Big)^{\frac{1}{q}},$
for $m\geq 1$. Roughly speaking, the Hörmander condition (1.3) corresponds to the $L_{1}$-integral regularity condition i.e. (1.5) with $q=1$, while (1.4) corresponds to the $L_{2}$-integral regularity condition, i.e., (1.5) with $q=2$. The fact $\delta_{1}(m)\lesssim_{d}\delta_{2}(m)$ shows that the $L_{2}$-integral regularity condition is stronger than the Hörmander condition (1.3), but weaker than the Lipschitz regularity condition (1.2). By applying the noncommutative Calderón-Zygmund decomposition in [Cadilhac2022Spectral], Hong, the first author, and Xu [hong2023maximal] established the weak type (1,1) boundedness for the noncommutative maximal truncated Calderón-Zygmund operators under the $L_{2}$-integral regularity condition. We refer the reader to [Cadilhac2022Spectral, Lai2024Noncommutative, Mei2009Pseudo-localization] for more results on weak (1,1) boundedness theory for noncommutative Calderón-Zygmund operators.
In this paper, we study a kernel smoothness condition that is slightly weaker than the $L_{q}$-integral regularity condition (1.5): the $L_{q}$-mean Hörmander condition ($1\leq q\leq\infty$), defined by
(1.6) $\sum_{m\geq 1}H_{q}(m)<\infty,$
where $H_{q}(m)$ is defined as
$\sup_{{\begin{subarray}y\in\operatorname{\mathbb{R}}^{d},\\ R>0\end{subarray}}}$ $\Big(\frac{(2^{m}R)^{d(q-1)}}{|B_{R}|}\int_{|v|\leq R}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x{\rm d}v\Big)^{\frac{1}{q}},$
for $m\geq 1$. Here, $B_{R}$ denotes the ball centered at the origin with radius $R>0$, and $|B_{R}|$ denotes its volume. It is not hard to see that $H_{q}(m)\lesssim_{d}\delta_{q}(m)$, meaning the $L_{q}$-mean Hörmander condition (1.6) is indeed weaker than the $L_{q}$-integral regularity condition (1.5). In what follows, we focus on the case $q=2$, i.e., the $L_{2}$-mean Hörmander condition:
(1.7) $\sum_{m\geq 1}H_{2}(m)<\infty,$
where $H_{2}(m)$ is defined as in (1.6) with $q=2$.
For two smoothness conditions $A$ and $B$, we write $A\prec B$ to mean that $A$ is weaker than $B$. Thus, we have the following conclusion:
[…]
This naturally leads to a question: can the $L_{2}$-mean Hörmander condition (1.7) still guarantee the noncommutative weak type $(1,1)$ boundedness of the maximal Calderón-Zygmund operators? This paper gives a positive answer to this question.
@@
Suppose $T$ is a Calderón-Zygmund operator defined as in (1.8) associated with a kernel satisfying (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}$ be defined as in (1.9). Suppose $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $1<p_{0}<\infty$, that is, for any $f\in L_{p_{0}}(\operatorname{\mathcal{N}})$,
@@
Our main strategy in this paper is to develop a refined noncommutative Calderón-Zygmund decomposition, adapted to kernels satisfying the $L_{2}$-mean Hörmander condition (1.7). The previous methods are insufficient in this context. On one hand, Parcet’s decomposition [Parcet2009Pseudo-localization] requires the stronger Lipschitz regularity condition (1.2) and may not be directly applied to the noncommutative maximal operator. On the other hand, while the decomposition by Cadilhac et al. [Cadilhac2022Spectral] can deal with the $L_{2}$-integral regularity condition (1.5) with $q=2$, it does not cover our target. To address this gap, we introduce a slightly new noncommutative Calderón-Zygmund decomposition achieved by further splitting of the bad function into a convolution part and a remainder part.
@@ 2. Noncommutative Calderón-Zygmund decomposition
The Calderón-Zygmund decomposition is a fundamental tool for studying the boundedness of singular integral operators. When transferring to the von Neumann algebraic setting, its noncommutative counterpart is primarily based on Cuculescu’s martingale projections [Cuculescu1971Martingales]. In this section, we first review Cuculescu’s construction, and then use it to develop our refined version of the noncommutative Calderón-Zygmund decomposition. Background on noncommutative $L_{p}$ spaces and maximal norms can be found in Appendix A.
@@ 2.1. Cuculescu’s construction
The following lemma, established in [Cuculescu1971Martingales] (see also [Parcet2009Pseudo-localization]), plays an important role in the noncommutative Calderón-Zygmund decomposition.
@@ 2.2. Noncommutative Calderón-Zygmund decomposition
We require the following property of this projection, whose proof can be found in [Cadilhac2022Spectral].
@@
Below, we construct a slightly new noncommutative Calderón-Zygmund decomposition inspired by [Cadilhac2022Spectral], and present its properties later.
@@
Define $h_{r}(x)=\frac{1}{r^{d}}h(\frac{x}{r})$ and $r_{n}=2^{-n-1}\sqrt{d}$. Then, $f$ admits the decomposition:
@@
$\widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{r_{n}},\quad\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}.$
@@
by using Lemma 2.2 again. Next, we insert a convolution operator to decompose $b$ further. Let $h_{r}$ and $r_{n}$ be defined as above, and let
$\left\{\begin{array}[]{ll}b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{r_{n}},\end{array}\right.\quad\text{and}\quad\left\{\begin{array}[]{ll}b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}.\end{array}\right$.
@@
Remark 2.5.
For any $n\in\operatorname{\mathbb{Z}}$ and $Q\in\operatorname{\mathcal{Q}}_{n}$, $r_{n}$ is the radius of the circumscribed ball of the cube $Q$.
@@
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}=0$ (i.e. the cancellation condition), and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$. Furthermore, $\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\leq 2\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
@@
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{d},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
@@
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{off},n}^{Q}=0$ (i.e. the cancellation condition), and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{off},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
@@
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{off},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
@@
The property $(i)$ follows from Parcet’s noncommutative Calderón-Zygmund decomposition, as detailed in [Cadilhac2018Weak] or [Cadilhac2022Spectral].
Consider $(ii)$. By the definition of $b_{\text{d},n}^{Q}$ in Theorem 2.4, we have $\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}=0$ for each $Q\in\operatorname{\mathcal{Q}}_{n}$. Note that $\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n}}=1$, we get
$\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}(x){\rm d}x=\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}(y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n}}(x-y){\rm d}x\Big){\rm d}y=0,$
which establishes the cancellation condition for $\widetilde{b}_{\text{d},n}^{Q}$.
We now prove the support condition. For each $Q\in\operatorname{\mathcal{Q}}$, from the construction of $\zeta_{s}$ in (2.3), we see
@@
Indeed, for a fixed $x\in(2s+1)Q$,
$\zeta_{s}(x)=1-\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}(x)\leq 1-p_{Q}.$
Hence,
@@
and the same holds for $p_{Q}\zeta_{s}(x)$, which lead to (2.4). By $\operatorname{supp}(h_{r_{n}})\subset B_{r_{n}}$,
$\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)$ $=\zeta_{s}(x)\bigg(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{r_{n}}(y)}b_{\text{d},n}^{Q}(u)h_{r_{n}}(y-u){\rm d}u\bigg)\zeta_{s}(x)$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{r_{n}}(y)\bigcap Q}\Big(\zeta_{s}(x)p_{Q}(f-f_{n})(u)p_{Q}\zeta_{s}(x)\Big)h_{r_{n}}(y-u){\rm d}u.$
Using (2.4), it suffices to show:
(2.5) $x\in(2s+1)Q_{u,n}\quad\text{if}\quad|x-y|\leq\Big(\frac{2s+1}{\sqrt{d}}-2\Big)r_{n}.$
In fact, taking $u\in B_{r_{n}}(y)\bigcap Q$, we have
$|x-c(Q_{u,n})|\leq|x-y|+|y-u|+|u-c(Q_{u,n})|\leq\frac{2s+1}{\sqrt{d}}r_{n},$
@@
The boundedness of $b_{\text{d},n}^{Q}$ results from the fact that $\phi(fp_{n})=\phi(f_{n}p_{n})$ and (2.2). This means,
@@
The proofs of (iii)(iii)$(v)$ follow by analogous arguments and are therefore omitted. ∎
@@
We will fix $s=[6\sqrt{d}]$ throughout this paper. Furthermore, the projection $\zeta_{s}$ is denoted simply by $\zeta$ when $s$ is fixed.

### v2

# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with a class of $L_{1}$-integral conditions

@@ Abstract.
We construct a slightly new noncommutative Calderón-Zygmund decomposition by further splitting the bad function. Using this tool, we prove the weak type (1,1) boundedness of noncommutative Calderón-Zygmund operators under a class of $L_{1}$-integral conditions, which are close to $L_{1}$-Dini conditions.
@@ 1. Introduction
Calderón-Zygmund operators, a well-known class of singular integral operators, were introduced by Calderón and Zygmund [3] in their seminal 1952 work. The kernel of such an operator satisfies specific conditions, including the size and regularity conditions, the latter of which was refined by Hörmander [8] in 1960.
@@
where $x$, $y$, $z\in\operatorname{\mathbb{R}}^{d}$. The Lipschitz regularity condition (1.2) can be relaxed to certain variants of smoothness conditions (see [6, 20, 21]). In the following, all Calderón-Zygmund kernels under consideration are of nonconvolution type.
Inspired by operator algebras, harmonic analysis, noncommutative geometry, and quantum probability (see [1, 4, 10, 11, 13, 14, 15, 16, 17]), noncommutative harmonic analysis has become an exciting field in recent years. However, generalizing the Calderón-Zygmund decomposition, a key technique in harmonic analysis first introduced in [3], to the noncommutative setting is challenging.
In 2009, Parcet [17] rigorously constructed a kind of noncommutative Calderón-Zygmund decomposition by using Cuculescu’s maximal weak type $(1,1)$ estimates for noncommutative martingales [5]. With this tool, Parcet established the weak type $(1,1)$ estimates for operator-valued Calderón-Zygmund singular integrals with the kernel satisfying the Lipschitz regularity condition (1.2) (Cadilhac later gave a shorter proof in [1]). Notably, the Hörmander condition
@@
which is weaker than (1.2), already guarantees the weak type $(1,1)$ boundedness in classical Calderón-Zygmund theory. It is therefore natural to ask whether the same still holds in the noncommutative setting? This problem remains open.
A step forward was made in 2022 by Cadilhac et al. [2]. They developed a more effective noncommutative Calderón-Zygmund decomposition without the off-diagonal term of the good function, which allowed them to prove the noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators if the kernel $K$ satisfies
@@
where $\ell(Q)$ and $c(Q)$ stand for the length and the center of the dyadic cube $Q$, respectively.
More generally, for $1\leq q\leq\infty$, we say that a kernel $K$ satisfies the $L_{q}$-Dini condition if
@@
where
$\delta_{q}(m):=\sup_{{\begin{subarray}{c}y\in\operatorname{\mathbb{R}}^{d},\ R>0\\ |v|<R\end{subarray}}}$ $\Big((2^{m}R)^{d(q-1)}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x\Big)^{\frac{1}{q}},$
for $m\geq 1$. The fact that $\delta_{1}(m)\lesssim_{d}\delta_{2}(m)$ shows that the $L_{2}$-Dini condition ((1.5) with $q=2$) is stronger than the $L_{1}$-Dini condition ((1.5) with $q=1$), but weaker than the Lipschitz regularity condition (1.2). Note that the $L_{1}$-Dini condition ((1.5) with $q=1$) corresponds to the Hörmander condition (1.3).
Using the noncommutative Calderón-Zygmund decomposition in [2], Hong, the first author, and Xu [7] established the weak type (1,1) boundedness for the noncommutative maximal truncated Calderón-Zygmund operators under the $L_{2}$-Dini condition ((1.5) with $q=2$). Further results on weak (1,1) boundedness theory for the noncommutative Calderón-Zygmund operators can be found in [2, 12, 14].
Since it remains open whether the $L_{1}$-Dini condition ((1.5) with $q=1$) guarantees the weak type $(1,1)$ boundedness, a natural intermediate problem is whether we can lower the exponent from $q=2$ to some $q$ satisfying $1\leq q<2$? This also appears to be quite challenging. In this paper, we focus on a class of $L_{1}$-integral conditions, which is close to the $L_{1}$-Dini condition ((1.5) with $q=1$). For a dyadic cube $Q$, we define the kernel average by
$\mathcal{K}_{Q}(x,y):=\frac{1}{|Q|}\int_{|v|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+v)-K(x,y)|{\rm d}v,$
and set
$\mathcal{H}(m):=\sup_{Q\ \text{dyadic}}\int_{2^{m}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{m+1}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}_{Q}(x,y){\rm d}x.$
We then say that the kernel satisfies the $L_{1}$-integral condition if
(1.6) $\sum_{m\geq 1}\mathcal{H}(m)<\infty.$
Compared with the $L_{1}$-Dini condition ((1.5) with $q=1$), the $L_{1}$-integral condition (1.6) replaces the supremum over $v$ by an average over the ball $\{v:|v|\leq\frac{\sqrt{d}}{2}\ell(Q)\}$, while the supremum over $y$ is taken inside the integral of $x$. The purpose of this paper is to show that this $L_{1}$-integral condition (1.6) is already sufficient for the weak type (1,1) boundedness of the noncommutative Calderón-Zygmund operators.
@@
Suppose $T$ is a Calderón-Zygmund operator defined as in (1.7) associated with a kernel satisfying (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}$ be defined as in (1.8). Suppose $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $1<p_{0}<\infty$, that is, for any $f\in L_{p_{0}}(\operatorname{\mathcal{N}})$,
@@
Our main strategy in this paper is to develop a refined noncommutative Calderón-Zygmund decomposition, along with related techniques, adapted to kernels satisfying the $L_{1}$-integral condition (1.6). Existing methods are inadequate for this purpose. Parcet’s decomposition [17] requires the stronger Lipschitz regularity condition (1.2) and may not be directly applied to the noncommutative maximal operators. Meanwhile, the decomposition by Cadilhac et al. [2] can only deal with the $L_{2}$-Dini condition ((1.5) with $q=2$) and therefore does not apply to our setting. To address this gap, we introduce a slightly new noncommutative Calderón-Zygmund decomposition achieved by further splitting of the bad function into a convolution part and a remainder part. In addition, we avoid using the Hölder inequality for the noncommutative square functions but instead employ more explicit factorizations for such terms.
@@ 2. Noncommutative Calderón-Zygmund decomposition
The Calderón-Zygmund decomposition is a fundamental tool for studying the boundedness of singular integral operators. When transferring to the von Neumann algebraic setting, its noncommutative counterpart is primarily based on Cuculescu’s martingale projections [5]. In this section, we first review Cuculescu’s construction, and then use it to develop our refined version of the noncommutative Calderón-Zygmund decomposition. Background on noncommutative $L_{p}$ spaces and maximal norms can be found in Appendix A.
@@ 2.1. Cuculescu’s construction
The following lemma, established in [5] (see also [17]), plays an important role in the noncommutative Calderón-Zygmund decomposition.
@@ 2.2. Noncommutative Calderón-Zygmund decomposition
We require the following property of this projection, whose proof can be found in [2].
@@
Below, we construct a slightly new noncommutative Calderón-Zygmund decomposition inspired by [2], and present its properties later.
@@
Define $h_{r}(x)=\frac{1}{r^{d}}h(\frac{x}{r})$. Then, $f$ admits the decomposition:
@@
$\widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)},\quad\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}.$
@@
by using Lemma 2.2 again. Next, we insert a convolution operator to decompose $b$ further. Let $h_{r}$ be defined as above. For a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, let
$\left\{\begin{array}[]{ll}b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)},\end{array}\right.\quad\text{and}\quad\left\{\begin{array}[]{ll}b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}.\end{array}\right$.
@@
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}=0$, and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$. Furthermore, $\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
@@
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{d},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
@@
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{off},n}^{Q}=0$, and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{off},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
@@
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{off},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
@@
$(i)$ follows directly from Parcet’s noncommutative Calderón-Zygmund decomposition, as detailed in [1] or [2].
Next, consider $(ii)$. The proof contains three parts: the cancellation condition, the support condition, and the boundedness.
By the definition of $b_{\text{d},n}^{Q}$ in Theorem 2.4, we have $\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}=0$ for each $Q\in\operatorname{\mathcal{Q}}_{n}$. Since $\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}=1$, it follows that
$\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}(x){\rm d}x=\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}(y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}(x-y){\rm d}x\Big){\rm d}y=0,$
which is the required cancellation property for $\widetilde{b}_{\text{d},n}^{Q}$.
Now we consider the support condition. For each $Q\in\operatorname{\mathcal{Q}}$, the construction of $\zeta_{s}$ in (2.3) implies
@@
Indeed, if $x\in(2s+1)Q$, then
$\zeta_{s}(x)=1-\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}(x)\leq 1-p_{Q},$
implying
@@
and the same holds for $p_{Q}\zeta_{s}(x)$, which lead to (2.4). Then, note that $\operatorname{supp}(h_{\frac{\sqrt{d}}{2}\ell(Q)})\subset B_{\frac{\sqrt{d}}{2}\ell(Q)}$, we write $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)$ as follows:
$\zeta_{s}(x)\bigg(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}}b_{\text{d},n}^{Q}(u)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}u\bigg)\zeta_{s}(x)$
$\quad=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{u\in\big(B_{\frac{\sqrt{d}}{2}\ell(Q)}(y)\bigcap Q\big)}\Big(\zeta_{s}(x)p_{Q}(f-f_{n})(u)p_{Q}\zeta_{s}(x)\Big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}u.$
As mentioned, we let $Q_{u,n}$ be the unique cube in $\operatorname{\mathcal{Q}}_{n}$ that contains $u$. Hence, with (2.4), it is enough to show that
(2.5) $x\in(2s+1)Q_{u,n}\quad\text{if}\quad|x-y|\leq(s-\sqrt{d})2^{-n}.$
In fact, for $u\in B_{\frac{\sqrt{d}}{2}\ell(Q_{u,n})}(y)\bigcap Q_{u,n}$,
$|x-c(Q_{u,n})|$ $\leq|x-y|+|y-u|+|u-c(Q_{u,n})|$
$\leq(s-\sqrt{d})\ell(Q_{u,n})+\frac{\sqrt{d}}{2}\ell(Q_{u,n})+\frac{\sqrt{d}}{2}\ell(Q_{u,n})$
$=s\ell(Q_{u,n})$
@@
The boundedness of $b_{\text{d},n}^{Q}$ results from the fact that $\phi(fp_{n})=\phi(f_{n}p_{n})$ together with (2.2). This means,
@@
$(iii)$ By the support condition in $(ii)$, we may restrict our attention to the nonconvolution part of $b_{\text{d},n}^{0}$. That is, by writing
$\zeta_{s}(x)b_{\text{d},n}^{0}\zeta_{s}(x)=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\zeta_{s}(x)b_{\text{d},n}^{Q}(y)\zeta_{s}(x)-\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\zeta_{s}(x)\widetilde{b}_{\text{d},n}^{Q}(y)\zeta_{s}(x),$
we only have to estimate the first sum, since the second one has already been considered in $(ii)$. The required estimate is obtained by using the same argument as in $(ii)$, and we therefore omit the details.
Finally, the proofs of $(iv)$ and $(v)$ follow from analogous reasoning and are also omitted.
@@
We will fix $s=[50\sqrt{d}]$ throughout this paper. Furthermore, the projection $\zeta_{s}$ is denoted simply by $\zeta$ when $s$ is fixed.